A motor gives gear A an angular acceleration of aA-udrad/sec2 where t is i seconds. f this gear is initially turning at ???-20 nd/s determine the angular a celeration of 9tul when t 2 sec. 20 rads 005 m 0.15 m
Answer #1

Solution :

We know that

\alpha _{A}r_{A}=\alpha _{B}r_{B}

\alpha _{A}=4t^{3}=4\times 2^{3}=32rad/s^{2}

32 × 0.05 0,05 OB 0.15 × 0.15

\alpha _{B}=10.67rad/s^{2}Answer